### v1

# On the quantum chromatic number of Hamming and generalized Hadamard graphs

Abstract
Quantum coloring finds applications in quantum cryptography and information. In this paper, we study the quantum chromatic numbers of Hamming graphs and a generalization of Hadamard graphs. We investigate the separation between the quantum and classical chromatic numbers of these graphs and determine the quantum chromatic numbers for some of them.
For the upper bounds of the quantum chromatic numbers, we develop a linear programming approach over the Hamming scheme to construct modulus-one orthogonal representations. For the lower bounds, we determine the minimum eigenvalues for some of these graphs to derive corresponding spectral lower bounds on their quantum chromatic numbers.
1 Introduction
Graph colouring plays a central role not only in combinatorics but also in quantum information theory and communication [2]. The quantum chromatic number of a graph $G$, denoted by $\chi_{Q}(G)$, was first suggested by Patrick Hayden (private communication, as reported in [2]) and independently introduced in [4]. Hadamard graphs $\varOmega_{n}$, which are defined when $n$ is a multiple of $4$, are graphs on vertex set consisting of all $\pm 1$}-vectors of length $n$, where two vertices are adjacent if and only if they are orthogonal. They can also be regarded as binary Hamming graphs with distance $n/2$, namely $H(n,2,n/2)$. These graphs provide a notable example of quantum advantage [2]: their quantum chromatic number satisfies $\chi_{Q}(\varOmega_{n})\leq n$, while the combinatorial result of Frankl and Rödl [9] implies that, for sufficiently large $n$, the classical chromatic number satisfies $\chi(\varOmega_{n})\geq(1+\varepsilon)^{n}$ for some $\varepsilon>0$, which yields an exponential separation between the quantum and classical chromatic numbers.
Despite its significance, few nontrivial lower bounds are known for the quantum chromatic number, and [10] showed that computing it is NP-hard in general. For a long time, apart from trivial classical graphs such as complete graphs, bipartite graphs, and cycles, the Hadamard graphs $\varOmega_{n}$ constituted the only known infinite family of graphs for which the quantum chromatic number could be determined.
Very recently, Cao, Feng, and Tan [5] determined the quantum chromatic number of another family of binary Hamming graphs, namely H(4t1,2,2t)H(4t-1,2,2t), which forms another known infinite family of graphs with an explicitly determined quantum chromatic number. In their work, they established an upper bound on $\chi_{Q}(H(n,2,d))$ for all $d\geq\frac{n}{2}$, while leaving the case $d<\frac{n}{2}$ as an open problem.
Motivated by [5], we study the quantum chromatic number and the corresponding separation properties of general $q$-ary Hamming graphs, as well as a natural generalization of Hadamard graphs.
The $q$-ary Hamming graph of length $n$ with distance $d$, denoted by $H(n,q,d)$, is the graph on the set of $n$-tuples over a $q$-ary alphabet, where two vertices are adjacent if and only if their coordinates differ in exactly $d$ positions. The collection of all such graphs forms a symmetric association scheme known as the Hamming scheme [6].
For Hamming graphs, we develop a linear programming method over the Hamming scheme to construct orthogonal representations, whose dimensions provide upper bounds on the quantum chromatic number. Using this method, we extend the upper bound on $\chi_{Q}(H(n,2,d))$ for $d\geq\tfrac{n}{2}$ obtained in [5] to the general case $H(n,q,d)$ with $d\geq\tfrac{(q-1)n}{q}$. Moreover, the versatility of our approach allows it to handle the case $d<\tfrac{(q-1)n}{q}$ as well, where we derive additional upper bounds that address the open question posed in [5]. Our main upper bounds are summarized as follows.
Theorem 1.1.
Let $n,q,d$ be positive integers with $q\geq 2$ and $d\leq n$.
  1. 1.
    If $d\geq\frac{(q-1)n}{q}$, then $\chi_{Q}(H(n,q,d))\leq qd$.
  2. 2.
    If $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, then $\chi_{Q}(H(n,q,d))\leq 2(q-1)^{2}\binom{n}{2}$.
  3. 3.
    If $d=\delta n$ for some $0<\delta<\frac{q-1}{q}$, then
    $\chi_{Q}(H(n,q,d))\leq q^{h_{q}\left(\frac{q-1-(q-2)\delta-2\sqrt{(q-1)\delta(1-\delta)}}{q}\right)n+o(n)},$
    where $h_{q}(x)=x\log_{q}(q-1)-x\log_{q}x-(1-x)\log_{1}(1-x)$ is the $q$-ary entropy function.
Remark 1.2.
Frankl and Rödl [9, Theorem 1.10] showed that for sufficiently large $n$ and all $d$ satisfying $\delta n<d<(1-\delta)n$ for some fixed $0<\delta<\tfrac{1}{2}$ and with $d$ even when $q=2$, there exists a constant $\varepsilon=\varepsilon(q,\delta)>0$ such that $\alpha(H(n,q,d))\leq(q-\varepsilon)^{n}$. Consequently, we have $\chi(H(n,q,d))\geq(\tfrac{q}{q-\varepsilon})^{n}$. Therefore, the first two cases in Theorem 1.1 yield an exponential separation between the quantum and classical chromatic numbers. The third case yields an MRRW-type upper bound; however, it does not lead to such a separation, which we leave as an open problem in Section 5.
In addition, we establish a Plotkin-type lower bound for the quantum chromatic number of Hamming graphs.
Theorem 1.3.
Let $n,q,d$ be positive integers with $q\geq 3$ and $d\leq n$.
  1. 1.
    If $d=\tfrac{(q-1)n}{q}$, then $\chi_{Q}(H(n,q,d))\geq(q-1)(n-1)+1$.
  2. 2.
    If $d\geq\frac{(q-1)n+1}{q}$, then $\chi_{Q}(H(n,q,d))\geq\dfrac{qd}{qd-(q-1)n}$.
Remark 1.4.
Combining the first case in Theorem 1.1, we obtain $(q-1)(n-1)+1\leq\chi_{Q}(H(n,q,\frac{(q-1)n}{q}))\leq(q-1)n$ and $\chi_{Q}(n,q,\frac{(q-1)n+1}{q})=(q-1)n+1$. There remains a gap of (q2)(q-2) between the upper and lower bounds of $\chi_{Q}(H(n,q,\tfrac{(q-1)n}{q}))$, and we leave bridging this gap as an open question in Section 5.
We consider the following generalization of the Hadamard graph with respect to an additive group $\mathbb{G}$, denoted by $\varOmega_{n}^{(\mathbb{G})}$, which is the graph on $\mathbb{G}^{n}$ where two vertices are adjacent if and only if each element of $\mathbb{G}$ appears exactly $n/|\mathbb{G}|$ times in their difference. Note that when $\mathbb{G}$ has order 2, so it must be the cyclic group $\mathbb{Z}_{2}$, $\varOmega_{n}^{(\mathbb{Z}_{2})}$ coincides with the ordinary Hadamard graph $\varOmega_{n}$. Moreover, this generalization naturally leads to a special association scheme. The composition of an $n$-tuple $x\in\mathbb{G}^{n}$, denoted by $\operatorname{comp}(x)$, is a $|\mathbb{G}|$-tuple of nonnegative integers $(d_{g}:g\in\mathbb{G})$ such that $\sum_{g\in\mathbb{G}}d_{g}=n$. (In this paper, we use bold font to denote a composition, e.g., $\boldsymbol{d}=(d_{g}:g\in\mathbb{G})$) The composition graph of length $n$ over $\mathbb{G}$ with composition $\boldsymbol{d}=(d_{g}:g\in\mathbb{G})$, denoted by $H_{C}(n,\mathbb{G},\boldsymbol{d})$, is the directed graph on $\mathbb{G}^{n}$ in which two vertices $x$ and $y$ form a directed edge from $x$ to $y$ if and only if $\operatorname{comp}(y-x)=\boldsymbol{d}$. Indeed, all such composition graphs form an asymmetric association scheme , which we refer to as the composition scheme.  [^1] To the best of our knowledge, this scheme was first introduced by Delsarte [6] in the case $\mathbb{G}=\mathbb{F}_{q}$, who referred to it as the spectral scheme, and was later further studied by Sookoo [13] in the same setting. Note that the composition scheme with respect to an order-$q$ additive group is a refinement of the $q$-ary Hamming scheme; in particular, when $q=2$, the composition scheme coincides with the binary Hamming scheme.
For generalized Hadamard graphs, we partially determine their quantum chromatic numbers in the following two cases:
Theorem 1.5.
Let $q$ be a positive integer and $n$ a positive integer divisible by $q$.
  1. 1.
    If $\frac{(q-1)n}{q}$ is even, then there exists $N=N(q)$ such that for every $n\geq N$,
    $\chi_{Q}(\varOmega_{n}^{(\mathbb{Z}_{q})})=n.$
  2. 2.
    If both $n$ and $q$ are prime powers, then
    $\chi_{Q}(\varOmega_{n}^{(\mathbb{F}_{q})})=n.$
The remaining part of this paper is organized as follows. In Section 2, we introduce some basics of graph theory and present fundamental facts on the quantum chromatic number. We further discuss Hamming and composition graphs and their eigenvalues. In Section 3, we prove Theorems 1.1 and 1.3 for the quantum chromatic number of Hamming graphs. In Section 4, we prove Theorem 1.5 for generalized Hadamard graphs. Finally, in Section 5, we conclude the paper and present several open problems for future research.
2 Preliminaries
All graphs considered in this paper are assumed to be simple, that is, undirected, with no loops or multiple edges. Directed graphs will be explicitly indicated when relevant.
Given two graphs $G$ and $H$, a homomorphism from $G$ to $H$ is a map $\phi:V(G)\to V(H)$ such that if $u$ and $v$ are adjacent in $G$, then $\phi(u)$ and $\phi(v)$ are adjacent in $H$. The chromatic number of $G$, denoted by $\chi(G)$, is the minimum positive integer $r$ such that there exists a homomorphism from $G$ to the complete graph on $r$ vertices. The clique number of $G$, denoted by $\omega(G)$, is the maximum positive integer $r$ such that there exists a homomorphism from the complete graph on $r$ vertices to $G$. The independence number of $G$, denoted by $\alpha(G)$, is the maximum size of an independent set in $G$ and is equal to the clique number of the complement of $G$. Observe that $\chi(G)\alpha(G)\geq|V(G)|$.
A quantum homomorphism from $G$ to $H$ is a set of Hermitian matrices
$\left\{P_{v,\alpha}:v\in V(G),\alpha\in V(H)\right\}\subseteq\mathbb{C}^{k\times k}$
for some positive integer $k$, satisfying:
  • (1)
    For every $v\in V(G)$, the set $\{P_{v,\alpha}:\alpha\in V(H)\}$ forms a complete orthogonal system, which means that $\sum_{\alpha\in V(H)}P_{v,\alpha}=I$ and $P_{v,\alpha}P_{v,\beta}=\delta_{\alpha,\beta}P_{v,\alpha}$ for all $\alpha,\beta\in V(H)$, where $\delta_{\alpha,\beta}$ equals 1 if $\alpha=\beta$ and 0 otherwise.
  • (2)
    For any two adjacent vertices $u,v\in V(G)$ and any two non-adjacent vertices $\alpha,\beta\in V(H)$, we have $P_{u,\alpha}P_{v,\beta}=O$.
Using quantum homomorphisms, the quantum chromatic number of $G$, denoted by $\chi_{Q}(G)$, is the minimum positive integer $r$ such that there exists a quantum homomorphism from $G$ to the complete graph on $r$ vertices. Clearly, any graph homomorphism naturally induces a quantum homomorphism. Therefore, we have $\chi_{Q}(G)\leq\chi(G)$.
We next recall two classical bounds for the quantum chromatic number: an upper bound based on modulus-one orthogonal representations and a spectral lower bound.
An orthogonal representation of a graph $G$ with dimension $K$ is a map $\rho:V(G)\to\mathbb{C}^{K}$ such that $\rho(u)$ and $\rho(v)$ are orthogonal with respect to the complex inner product for all adjacent vertices $u,v\in V(G)$. Moreover, the representation $\rho$ is called modulus-one if all coordinates of $\rho(v)$ have modulus one for every $v\in V(G)$. Let $\xi(G)$ and $\xi^{\prime}(G)$ denote the minimum dimension of an orthogonal representation and of a modulus-one orthogonal representation, respectively.
As established by Cameron et al. in [4], the quantum chromatic number of $G$ satisfies the following upper bound:
Lemma 2.1.
$\chi_{Q}(G)\leq\xi^{\prime}(G)$.
On the other hand, spectral techniques provide a lower bound for $\chi_{Q}(G)$. Let $G$ be a graph on $n$ vertices, and let $\lambda_{1}\geq\dots\geq\lambda_{n}$ denote the eigenvalues of its adjacency matrix. The following Hoffman-type bound was established in [8]:
Lemma 2.2.
$\chi_{Q}(G)\geq 1-\frac{\lambda_{1}}{\lambda_{n}}$.
It is well known that if $G$ is an $r$-regular graph, then $\lambda_{1}=r$ is the largest eigenvalue in absolute value. Note that both Hamming graphs and generalized Hadamard graphs are regular. More precisely, $H(n,q,d)$ is $(q-1)^{d}\binom{n}{d}-$regular, and $\varOmega_{n}^{(\mathbb{G})}$ is $\binom{n}{n/q,\dots,n/q}-$regular.
The Cayley graph on a group $\mathbb{G}$ with a generating set $S\subseteq\mathbb{G}$, denoted by $\operatorname{Cay}(\mathbb{G},S)$, is the directed graph on vertex set $\mathbb{G}$ where two vertices $g$ and $h$ form a directed edge $(g,h)$ if and only if $g^{-1}h\in S$. In particular, if $S^{-1}=S$, then $\operatorname{Cay}(\mathbb{G},S)$ is an undirected graph.
In fact, Hamming and composition graphs can be represented as Cayley graphs. Given a $q$-ary alphabet $\varSigma$, equip it with an arbitrary additive group structure $\mathbb{G}=(\varSigma,+)$. For any $n$-tuple $x\in\mathbb{G}^{n}$, define its Hamming weight $w_{\mathrm{H}}(x)$ as the number of coordinates not equal to $0_{\mathbb{G}}$, and its composition $\operatorname{comp}(x)=\boldsymbol{i}=(i_{g}:g\in\mathbb{G})$, where $i_{g}$ is the number of coordinates of $x$ equal to $g$. The Hamming graph $H(n,q,i)$ and the composition graph $H_{C}(n,\mathbb{G},\boldsymbol{i})$ can be represented as the Cayley graphs $\operatorname{Cay}(\mathbb{G}^{n},S_{i})$ and $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{i}})$, respectively, where $S_{i}=\{x\in\mathbb{G}^{n}:w_{\mathrm{H}}(x)=i\}$ and $S_{\boldsymbol{i}}=\{x\in\mathbb{G}^{n}:\operatorname{comp}(x)=\boldsymbol{i}\}$. In particular, when $q\mid n$, we simply write the composition $(n/q,\dots,n/q)$ in bold as $\boldsymbol{n/q}$, and the generalized Hadamard graph $\varOmega_{n}^{(\mathbb{G})}=H_{C}(n,\mathbb{G},\boldsymbol{n/q})$ can be represented as the Cayley graph $\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{n/q}})$.
Note that both Hamming and Composition graphs are Cayley graphs over an Abelian group. The eigenvalues of an Abelian Cayley graph can be expressed in an elegant way. Before presenting their eigenvalues, we first recall some basics of the characters of an Abelian group.
Let $\mathbb{G}$ be a finite Abelian group. A character $\varphi$ of $\mathbb{G}$ is a homomorphism from $\mathbb{G}$ to the multiplicative group of complex numbers, i.e., $\varphi:\mathbb{G}\to\mathbb{C}^{\times}$, satisfying $\varphi(xy)=\varphi(x)\varphi(y)$. Let $\widehat{\mathbb{G}}$ denote the set of all characters of $\mathbb{G}$. For each $x\in\mathbb{G}$, we have $\varphi(x)^{|\mathbb{G}|}=\varphi(|\mathbb{G}|x)=\varphi(0_{\mathbb{G}})=1$. Thus the values of $\varphi$ are $|\mathbb{G}|$-th roots of unity. Moreover, $\widehat{\mathbb{G}}$ forms an Abelian group under entrywise multiplication, and there is a canonical isomorphism from $\mathbb{G}$ to $\widehat{\mathbb{G}}$. Under this isomorphism, we can relabel $\widehat{\mathbb{G}}=\{\varphi_{g}:g\in\mathbb{G}\}$ such that $\varphi_{g+h}=\varphi_{g}\varphi_{h}$ and $\varphi_{g}(h)=\varphi_{h}(g)$ for all $g,h\in\mathbb{G}$.
It is well known that for $\mathbb{G}=\mathbb{Z}_{q}^{n}$, the character group is given by $\widehat{\mathbb{Z}_{q}^{n}}=\{\varphi_{x}:x\in\mathbb{Z}_{q}^{n}\}$, with $\varphi_{x}(y)=\zeta_{q}^{x\cdot y}$ for all $y\in\mathbb{Z}_{q}^{n}$. Similarly, for $\mathbb{G}=\mathbb{F}_{q}^{n}$, where $q$ is a power of a prime $p$, we have $\widehat{\mathbb{F}_{q}^{n}}=\{\psi_{\alpha}:\alpha\in\mathbb{F}_{q}^{n}\}$, and the characters are defined by $\psi_{\alpha}(\beta)=\zeta_{p}^{\operatorname{Tr}(\alpha\cdot\beta)}$ for all $\beta\in\mathbb{F}_{q}^{n}$, where $\operatorname{Tr}(\cdot)$ denote the trace function from $\mathbb{F}_{q}$ to $\mathbb{F}_{p}$. In both cases, the dot product $x\cdot y$ is defined as $\sum_{i=1}^{n}x_{i}y_{i}$, and $\zeta_{r}$ be the primitive $r$-th roots of unity.
We now recall a classical result on the spectrum of Abelian Cayley graphs using characters.
Lemma 2.3 ([12]).
Let $\mathbb{G}$ be a finite Abelian group, and let $S$ be a subset of $\mathbb{G}$. Then the eigenvalues of the Cayley graph $\operatorname{Cay}(\mathbb{G},S)$ are given by
$\lambda_{g}=\sum_{s\in S}\varphi_{g}(s),\quad g\in{\mathbb{G}}.$
Moreover, the eigenvector corresponding to $\lambda_{\varphi}$ is $(\varphi(x):x\in\mathbb{G})^{\top}$.
As a consequence, this result provides an explicit expression for the eigenvalues of Hamming and composition graphs.
For Hamming graphs, the eigenvalues of $H(n,q,i)\cong\operatorname{Cay}(\mathbb{G}^{n},S_{i})$ are given by
$\lambda_{x}=\sum_{y\in S_{i}}\varphi_{x}(y),\quad x\in\mathbb{G}^{n}.$
A direct calculation in [6] shows that $\lambda_{x}$ depends only on the Hamming weight of $x$. More precisely, if $w_{\mathrm{H}}(x)=j$, then $\lambda_{x}$ equals the degree-$i$ Krawtchouk polynomial evaluated at $j$, denoted by $K_{i}(j)$, which is defined by
$K_{i}(j)=\sum_{k=0}^{i}(-1)^{k}(q-1)^{i-k}\binom{j}{k}\binom{n-j}{i-k}$.
This fact leads naturally to the following reciprocal property, which can be derived by a simple double counting argument:
$(q-1)^{j}\binom{n}{j}K_{i}(j)=\sum_{x\in S_{j}}\sum_{y\in S_{i}}\varphi_{x}(y)=\sum_{y\in S_{i}}\sum_{x\in S_{j}}\varphi_{y}(x)=(q-1)^{i}\binom{n}{i}K_{j}(i).$
For composition graphs, the eigenvalues of $H(n,\mathbb{G},\boldsymbol{i})\cong\operatorname{Cay}(\mathbb{G}^{n},S_{\boldsymbol{i}})$ are given by
$\mu_{x}=\sum_{y\in S_{\boldsymbol{i}}}\varphi_{x}(y),\quad x\in\mathbb{G}^{n}.$
Let $\boldsymbol{z}=(z_{g}:g\in\mathbb{G})$ be a tuple of indeterminates, and denote the monomial $\prod_{g\in\mathbb{G}}z_{g}^{i_{g}}$ by $\boldsymbol{z}^{\boldsymbol{i}}$. We use the bracket notation $\cdot$] to denote the coefficient extraction operator; that is, $\boldsymbol{z}^{\boldsymbol{i}}]f(\boldsymbol{z})$ gives the coefficient of $\boldsymbol{z}^{\boldsymbol{i}}$ in the expansion of $f(\boldsymbol{z})$.
To encode the above character sums over $n$-tuples in $S_{\boldsymbol{i}}$, we consider the generating function
$\prod_{k=1}^{n}\left(\sum_{g\in\mathbb{G}}\varphi_{x_{k}}(g)z_{g}\right),$
where for each $k$, $\varphi_{x_{k}}$ is the character of $\mathbb{G}$ corresponding to the $k$-th coordinate of $x$.
Expanding the product, each monomial corresponds to an $n$-tuple $y\in\mathbb{G}^{n}$, with the exponent of $z_{g}$ equal to the number of coordinates of $y$ that are equal to $g$. It follows that the coefficient of $\boldsymbol{z}^{\boldsymbol{i}}$ in this expansion is precisely the sum of the product of characters over all $y\in S_{\boldsymbol{i}}$, that is,
$\sum_{y\in S_{\boldsymbol{i}}}\varphi_{x}(y)=\sum_{y\in S_{\boldsymbol{i}}}\prod_{k=1}^{n}\varphi_{x_{k}}(y_{k})=[\boldsymbol{z}^{\boldsymbol{i}}]\prod_{k=1}^{n}\left(\sum_{g\in\mathbb{G}}\varphi_{x_{k}}(g)z_{g}\right).$
From the second expression of $\mu_{x}$, it is clear that it depends only on the composition of $x$. That is, if $\mathrm{comp}(x)=\boldsymbol{j}=(j_{g}:g\in\mathbb{G})$, then
$\mu_{x}=[\boldsymbol{z}^{\boldsymbol{i}}]\prod_{k=1}^{n}\left(\sum_{g\in\mathbb{G}}\varphi_{x_{k}}(g)z_{g}\right)=[\boldsymbol{z}^{\boldsymbol{i}}]\prod_{h\in\mathbb{G}}\left(\sum_{g\in\mathbb{G}}\varphi_{h}(g)z_{g}\right)^{j_{h}}.$
We define the generalized Krawtchouk polynomial with respect to $\mathbb{G}$  [^2] In the special case $\mathbb{G}=\mathbb{F}_{q}$, this polynomial coincides with the one studied by Sookoo [13]. as
$K^{(\mathbb{G})}_{\boldsymbol{i}}(\boldsymbol{j})=[\boldsymbol{z}^{\boldsymbol{i}}]\prod_{h\in\mathbb{G}}\left(\sum_{g\in\mathbb{G}}\varphi_{h}(g)z_{g}\right)^{j_{h}},$
Finally, by a simple double-counting argument, the generalized Krawtchouk polynomials also satisfy the reciprocal law:
$\binom{n}{\boldsymbol{j}}K^{(\mathbb{G})}_{\boldsymbol{i}}(\boldsymbol{j})=\sum_{x\in S_{\boldsymbol{j}}}\sum_{y\in S_{\boldsymbol{i}}}\varphi_{x}(y)=\sum_{y\in S_{\boldsymbol{i}}}\sum_{x\in S_{\boldsymbol{j}}}\varphi_{y}(x)=\binom{n}{\boldsymbol{i}}K^{(\mathbb{G})}_{\boldsymbol{j}}(\boldsymbol{i}).$
where we use the notation $\binom{n}{\boldsymbol{i}}$ for the multinomial coefficient $\frac{n!}{\prod_{g\in\mathbb{G}}i_{g}!}$.
3 Quantum chromatic number of some Hamming graphs
The goal of this section is to present the proofs of Theorems 1.1 and 1.3.
3.1 A linear programming approach to orthogonal representations
In this subsection, we derive upper bounds on $\xi^{\prime}(H(n,q,d))$ using a linear programming approach. We begin by recalling some basic facts about the Hamming scheme.
Let $A_{i}$ be the adjacency matrix of $H(n,q,i)$ for $i=0,1,\dots,n$. Then $A_{0}=I,A_{1},\dots,A_{n}$ span the Bose–Mesner algebra of the Hamming scheme, denoted by
$\mathbb{A}=\operatorname{span}_{\mathbb{C}}\{A_{0},A_{1},\ldots,A_{n}\}.$
Denote $|\varphi_{x}\rangle=(\varphi_{x}(y):y\in\mathbb{G}^{n})^{\top}$. As shown in the previous section, we have
$A_{i}\cdot|\varphi_{x}\rangle=K_{i}(w_{\mathrm{H}}(x))\cdot|\varphi_{x}\rangle,\quad i=0,1,\dots,n.$
This implies that all matrices in $\mathbb{A}$ can be simultaneously diagonalized. Consequently, viewing $\mathbb{C}^{q^{n}}$ as an $\mathbb{A}$-module, it decomposes orthogonally into a direct sum of $\mathbb{A}$-submodules:
$\mathbb{C}^{q^{n}}=V_{0}\perp V_{1}\perp\dots\perp V_{n},$
where $V_{j}=\operatorname{span}_{\mathbb{C}}\{|\varphi_{x}\rangle:w_{\mathrm{H}}(x)=j\}$. Let $E_{j}$ be the orthogonal projection from $\mathbb{C}^{q^{n}}$ onto $V_{j}$, namely
$E_{j}=\sum_{x\in S_{j}}|\varphi_{x}\rangle\langle\varphi_{x}|,\quad j=0,1,\dots,n,$
where $\langle\varphi_{x}|$ denote the conjugate transpose of $|\varphi_{x}\rangle$. Then $E_{0}=J,E_{1},\dots,E_{n}$ form a complete orthogonal system, and each adjacency matrix $A_{i}$ can be expressed as a linear combination of these projectors, with the corresponding eigenvalues as coefficients:
$A_{i}=K_{i}(0)E_{0}+K_{i}(1)E_{1}+\dots+K_{i}(n)E_{n},\quad i=0,1,\dots,n.$
Therefore, $\mathbb{A}$ lies in the span of $E_{0},E_{1},\dots,E_{n}$. Since $A_{0},A_{1},\dots,A_{n}$ are clearly linearly independent, we have $\dim(\mathbb{A})=n+1$. By comparing dimensions, it follows that the span of $E_{0},E_{1},\dots,E_{n}$ is exactly $\mathbb{A}$. It is well known (see, e.g., [6]) that the following change-of-basis formula holds:
$E_{j}=\frac{1}{q^{n}}\bigl(K_{j}(0)A_{0}+K_{j}(1)A_{1}+\cdots+K_{j}(n)A_{n}\bigr),\quad j=0,1,\dots,n.$
The following lemma presents an explicit construction of a modulus-one orthogonal representation of Hamming graphs.
Lemma 3.1.
The quantity $\xi’(H(n,q,d))$ is upper bounded by the value of any feasible solution to the following linear program:
[…]
Proof.
Let $(c_{0},\dots,c_{n})$ be a feasible solution to the linear program, and define
$M=\sum_{i=0}^{n}c_{i}E_{i}.$
Observe that each $E_{i}$ admits a decomposition
$E_{i}=\varPhi_{i}\varPhi_{i}^{\dagger},$
where the columns of $\varPhi_{i}$ are the vectors $|\varphi_{x}\rangle$ for $x\in S_{i}$, and the dagger $\cdot^{\dagger}$ denotes the conjugate transpose operator. It follows that
$M=\sum_{i=0}^{n}c_{i}\varPhi_{i}\varPhi_{i}^{\dagger}.$
Using the coefficients $c_{i}$, construct a matrix $N$ by concatenating $c_{i}$ copies of $\varPhi_{i}$ side by side for each $i$, namely,
$N=\Big(\underbrace{\varPhi_{0},\dots,\varPhi_{0}}_{\text{$c_{0}$ copies}},\underbrace{\varPhi_{1},\dots,\varPhi_{1}}_{\text{$c_{1}$ copies}},\dots,\underbrace{\varPhi_{n},\dots,\varPhi_{n}}_{\text{$c_{n}$ copies}}\Big).$
Note that each $c_{i}$ is a natural number and that their sum $\sum_{i=0}^{n}c_{i}$ is positive, hence $N$ is not empty. Let $k=\sum_{i=0}^{n}(q-1)^{i}\binom{n}{i}c_{i}$ be the number of columns of $N$, and define a map $\rho:\mathbb{G}^{n}\to\mathbb{C}^{k}$ that sends each vertex $x\in\mathbb{G}^{n}$ to the row of $N$ indexed by $x$. By construction, the complex inner product of $\rho(x)$ and $\rho(y)$ coincides with the $(x,y)$-entry of $M$, i.e.,
$\langle\rho(x),\rho(y)\rangle=M_{x,y}.$
On the other hand, applying the change-of-basis formula
$E_{j}=\frac{1}{q^{n}}\sum_{i=0}^{n}K_{j}(i)A_{i},$
we can rewrite $M$ in terms of the adjacency matrices:
$M=\sum_{i=0}^{n}c_{i}E_{i}=\sum_{i=0}^{n}c_{i}\left(\frac{1}{q^{n}}\sum_{j=0}^{n}K_{i}(j)A_{j}\right)=\frac{1}{q^{n}}\sum_{j=0}^{n}\left(\sum_{i=0}^{n}c_{i}K_{i}(j)\right)A_{j}.$
By feasibility, the coefficient of $A_{d}$ in this expansion is zero. Hence, for any adjacent vertices $x,y\in H(n,q,d)$,
$\langle\rho(x),\rho(y)\rangle=M_{x,y}=0,$
showing that $\rho$ is indeed an orthogonal representation of dimension $\sum_{i=0}^{n}(q-1)^{i}\binom{n}{i}c_{i}$, and is clearly modulus-one. ∎
Corollary 3.2.
If $d\geq\frac{(q-1)n}{q}$, then $\xi^{\prime}(H(n,q,d))\leq qd$.
Proof.
Since $d\geq\frac{(q-1)n}{q}$, we have $K_{1}(d)=(q-1)n-qd\leq 0$. Therefore, by setting $c_{0}=-K_{1}(d)$, $c_{1}=1$, and $c_{i}=0$ for all other $i$, we obtain a feasible solution. Consequently, we have
$\xi^{\prime}(H(n,q,d))\leq-K_{1}(d)+(q-1)\binom{n}{1}=qd-(q-1)n+(q-1)n=qd.$
Corollary 3.3.
If $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, then $\xi^{\prime}(H(n,q,d))\leq 2(q-1)^{2}\binom{n}{2}$.
Proof.
Observe that
$K_{2}(d)=\frac{q^{2}}{2}d^{2}-\left(q(q-1)n-\frac{q(q-2)}{2}\right)d+\frac{(q-1)^{2}n(n-1)}{2},$
which is negative if and only if $\frac{(q-1)n}{q}-\sqrt{\frac{(q-1)n}{q^{2}}+\frac{(q-2)^{2}}{4}}-\frac{q-2}{2q}<d<\frac{(q-1)n}{q}+\sqrt{\frac{(q-1)n}{q^{2}}+\frac{(q-2)^{2}}{4}}-\frac{q-2}{2q}$. Hence, when $\frac{(q-1)n}{q}-\frac{\sqrt{(q-1)n}}{q}<d<\frac{(q-1)n}{q}$, we have $K_{2}(d)<0$. Therefore, setting $c_{0}=-K_{2}(d)$, $c_{2}=1$, and $c_{i}=0$ for all other $i$ yields a feasible solution.
It follows that
$\xi^{\prime}(H(n,q,d))\leq-K_{2}(d)+(q-1)^{2}\binom{n}{2}\leq 2(q-1)^{2}\binom{n}{2}$,
where the last inequality holds since $(q-1)^{2}\binom{n}{2}$ is the maximum eigenvalue in absolute value of $H(n,q,2)$ and $K_{2}(d)$ is another eigenvalue of this graph. ∎
Corollary 3.4.
If $d=\delta n$ for some $0<\delta<\frac{q-1}{q}$, then for sufficiently large $n$, we have
$\xi^{\prime}(H(n,q,d))\leq q^{h_{q}\left(\frac{q-1-(q-2)\delta-2\sqrt{(q-1)\delta(1-\delta)}}{q}\right)n+o(n)}.$
Proof.
It is well known (see, e.g., [11]) that the degree $d$ Krawtchouk polynomial $K_{d}(z)$ has $d$ distinct real zeros $z_{1}^{(d)}<\dots<z_{d}^{(d)}$, and that each interval between two consecutive zeros contains at least one integer. Moreover, $z_{1}^{(d)}$ is asymptotically given by
$\lim_{n\to\infty}\frac{z_{1}^{(\delta n)}}{n}=\frac{q-1-(q-2)\delta-2\sqrt{(q-1)\delta(1-\delta)}}{q}.$
Let $\lceil z_{1}^{(d)}\rceil$ denote the smallest integer greater than or equal to $z_{1}^{(d)}$. From the properties listed above, we know that $\lceil z_{1}^{(d)}\rceil\in(z_{1}^{(d)},z_{2}^{(d)})$. Since $K_{d}(0)=(q-1)^{d}\binom{n}{d}>0$, by the definition of $\lceil z_{1}^{(d)}\rceil$ and the continuity of $K_{d}(z)$, it follows that $K_{d}(\lceil z_{1}^{(d)}\rceil)\leq 0$. Hence, by the reciprocal law, we have $K_{\lceil z_{1}^{(d)}\rceil}(d)\leq 0$.
Therefore, setting $c_{0}=-K_{\lceil z_{1}^{(d)}\rceil}(d)$, $c_{\lceil z_{1}^{(d)}\rceil}=1$, and $c_{i}=0$ for all other $i$ yields a feasible solution. Consequently, we have
$\begin{split}\xi^{\prime}(H(n,q,d))&\leq-K_{\lceil z_{1}^{(d)}\rceil}(d)+(q-1)^{\lceil z_{1}^{(d)}\rceil}\binom{n}{\lceil z_{1}^{(d)}\rceil}\\ &\leq 2(q-1)^{\lceil z_{1}^{(d)}\rceil}\binom{n}{\lceil z_{1}^{(d)}\rceil}\\ &\leq q^{h_{q}\left(\frac{q-1-(q-2)\delta-2\sqrt{(q-1)\delta(1-\delta)}}{q}\right)n+o(n)},\end{split}$
where the second inequality holds since $(q-1)^{\lceil z_{1}^{(d)}\rceil}\binom{n}{\lceil z_{1}^{(d)}\rceil}$ is the maximum eigenvalue in absolute value of $H(n,q,\lceil z_{1}^{(d)}\rceil)$ while $K_{\lceil z_{1}^{(d)}\rceil}(d)$ is another eigenvalue of this graph; the last inequality follows from the well-known entropy estimate for Hamming spheres, namely, $(q-1)^{t}\binom{n}{t}\leq q^{h_{q}(\frac{t}{n})n}$. ∎
Proof of Theorem 1.1.
Theorem 1.1 follows directly from Lemma 2.1 and the combination of Corollaries 3.2, 3.3 and 3.4. ∎
3.2 Minimum eigenvalue of certain Hamming graphs
In this subsection, we derive a lower bound for the quantum chromatic number of $H(n,q,d)$ using Lemma 2.2. Since the Hamming graph $H(n,q,d)$ is $(q-1)^{d}\binom{n}{d}-$regular, its maximum eigenvalue equals $(q-1)^{d}\binom{n}{d}$. Consequently, we only need to focus on the minimum eigenvalue of $H(n,q,d)$.
The minimum eigenvalue is an important quantity relevant to many combinatorial problems, such as the maximum cut and intersecting families. However, determining it is generally a challenging task. Motivated by semidefinite programming approaches to the max-cut problem on Hamming graphs, Van Dam and Sotirov [14] conjectured that for $d\geq\frac{(q-1)n+1}{q}$, with $d$ even when $q=2$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)$. Alon and Sudakov [1] proved this for $q=2$ with $n$ large and $d$ fixed. Dumer and Kapralova [7] proved it for $q=2$ and all $n$. Finally, this conjecture was proved by Brouwer et al. [3]. We formulate it as follows:
Lemma 3.5 ([3, Theorem 1.4]).
For $d\geq\frac{(q-1)n+1}{q}$ with $q\geq 3$, the minimum eigenvalue of $H(n,q,d)$ is $K_{d}(1)=-(q-1)^{d}\binom{n}{d}\frac{qd-(q-1)n}{(q-1)n}$.
In the remaining part of this subsection, we focus on the minimum eigenvalue of $H(n,q,d)$ for $d=\frac{(q-1)n}{q}$ with $q\geq 3$. First, we have the following observation.
Lemma 3.6.
Let $G$ be an $r$-regular graph. Let $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ be the eigenvalues of the adjacency matrix of $G$, with corresponding multiplicities $m_{1},m_{2},\dots,m_{k}$. Then, for each $i=1,2,\dots,k$, we have
$\left(\frac{\lambda_{i}}{r}\right)^{2}\leq\frac{|V(G)|}{rm_{i}}.$
Proof.
Let $A$ be the adjacency matrix of the $r$-regular graph $G$. Since each diagonal entry of $A^{2}$ equals $r$, we have $\mathrm{tr}(A^{2})=|V(G)|r$. On the other hand, as the eigenvalues of $A^{2}$ are $\lambda_{i}^{2}$ with multiplicities $m_{i}$, we have $\mathrm{tr}(A^{2})=\sum_{i=1}^{k}m_{i}\lambda_{i}^{2}$. Equating the two expressions gives $\sum_{i}m_{i}\lambda_{i}^{2}=|V(G)|r$. Hence $m_{i}\lambda_{i}^{2}\leq|V(G)|r$ for each $i$, and dividing both sides by $m_{i}r^{2}$ yields $\left(\frac{\lambda_{i}}{r}\right)^{2}\leq\frac{|V(G)|}{rm_{i}}$, as claimed. ∎
Lemma 3.7.
For $d=\frac{(q-1)n}{q}$ with $q\geq 3$, the minimum eigenvalue of Hamming graph $H(n,q,d)$ is $K_{d}(2)=-(q-1)^{d}\binom{n}{d}\frac{1}{(q-1)(n-1)}$.
Proof.
We only consider the following cases: (i) $q\geq 5$ and $n\geq 5$, (ii) $q=4$ and $n\geq 8$, and (iii) $q=3$ and $n\geq 18$, since the remaining finitely many cases can be checked directly.
As discussed in Section 2, the eigenvalues of $H(n,q,d)$ are given by $K_{d}(i)$ for $i=0,1,\dots,n$, with (formal) multiplicities $(q-1)^{i}\binom{n}{i}$. By Lemma 3.6, we have
$\left(\frac{K_{d}(i)}{(q-1)^{d}\binom{n}{d}}\right)^{2}\leq\frac{q^{n}}{(q-1)^{d}\binom{n}{d}(q-1)^{i}\binom{n}{i}}$.
To show that $K_{d}(2)$ is indeed the minimum eigenvalue, we observe that $K_{d}(0)$ attains the maximum eigenvalue while $K_{d}(1)=0$. Therefore, it suffices to show that for each $3\leq i\leq n$,
$\frac{q^{n}}{(q-1)^{d}\binom{n}{d}(q-1)^{i}\binom{n}{i}}\leq\left(\frac{K_{d}(2)}{(q-1)^{d}\binom{n}{d}}\right)^{2}.$
Since $K_{d}(2)=-(q-1)^{d}\binom{n}{d}\frac{1}{(q-1)(n-1)}$, after rearranging the inequality it suffices to show that
$(q-1)^{i}\binom{n}{i}\geq\frac{q^{n}(q-1)^{2}(n-1)^{2}}{(q-1)^{d}\binom{n}{d}}$.
The quantity $(q-1)^{i}\binom{n}{i}$ is unimodal in $i$ and attains its maximum at $i=\frac{(q-1)n}{q}$. Therefore,
$(q-1)^{i}\binom{n}{i}\geq\min\!\left\{(q-1)^{3}\binom{n}{3},\,(q-1)^{n}\right\}\geq(q-1)^{3}\binom{n}{3}$,
where the last inequality holds for all three cases (i)–(iii).
Let
$R(q,n)=\frac{(q-1)^{3}\binom{n}{3}(q-1)^{d}\binom{n}{d}}{q^{n}(q-1)^{2}(n-1)^{2}}.$
Using Stirling’s approximation
$\sqrt{2\pi n}\frac{n^{n}}{e^{n}}\leq n!\leq\sqrt{2\pi n}\frac{n^{n}}{e^{n}}e^{\frac{1}{12n}},$
we obtain
$\begin{split}R(q,n)&\geq\frac{(n-2)(q-1)}{6q^{n}}(q-1)^{d}\binom{n}{d}\\ &\geq\frac{(n-2)(q-1)}{6q^{n}}\cdot q^{n}\frac{q}{\sqrt{2\pi n(q-1)}}e^{-\frac{q^{2}}{12(q-1)n}}\\ &\gtrsim_{q}\sqrt{n}.\end{split}$
Hence $R(q,n)$ grows at least on the order of $\sqrt{n}$ and is monotone increasing for large $n$. Consequently, for sufficiently large $n$, we have $R(q,n)\geq 1$. A direct computation further confirms that this already holds at the boundary values: $R(5,5)\geq 1.023\geq 1$, $R(4,8)\geq 1.067\geq 1$, and $R(3,18)\geq 1.108\geq 1$. This completes the proof. ∎
Proof of Theorem 1.3.
Together with Lemmas 2.2, 3.5 and 3.7, this yields the proof. ∎
4 Quantum chromatic number of generalized Hadamard graphs
In this section, we discuss the quantum chromatic number of the generalized Hadamard graph $\varOmega_{n}^{(\mathbb{G})}$. These graphs admit a natural modules-one orthogonal representation of dimension $n$, which we briefly discuss in the following lemma.
Lemma 4.1.
$\xi^{\prime}(\varOmega_{n}^{(\mathbb{G})})\leq n$.
Proof.
Let $\varphi$ be a non-trival character of $\mathbb{G}$, consider the map $\rho:\mathbb{G}^{n}\to\mathbb{C}^{n}$ that send each vertex $x$ to the vector $(\varphi(x_{1}),\dots,\varphi(x_{n}))$, therefore for any two adjacent vertices $x,y$, i.e., $\operatorname{comp}(x-y)=\boldsymbol{n/q}$, we have
$\langle\rho(x),\rho(y)\rangle=\sum_{k=1}^{n}\overline{\varphi(x_{k})}{\varphi(y_{k})}=\sum_{k=1}\varphi(y_{k}-x_{k})=\frac{n}{q}\sum_{g\in\mathbb{G}}\varphi(g)=0,$
where the last equality follows from the first orthogonality relation of characters. It is clear that the orthogonal representation $\rho$ is modulus-one, since the values of $\varphi$ are $|\mathbb{G}|$-th roots of unity. ∎
In the remaining part of this section, we focus on the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{G})}$, considering the cases $\mathbb{G}=\mathbb{Z}_{q}$ in Section 4.1 and $\mathbb{G}=\mathbb{F}_{q}$ in Section 4.2.
4.1 Minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$
Here, we apply the same approach as in Section 3.2 to determine the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$.
We begin by establishing further properties of the generalized Krawtchouk polynomial $K^{(\mathbb{Z}_{q})}_{\boldsymbol{i}}(\boldsymbol{j})$.
From the discussion in Section 2, the eigenvalues of $\varOmega_{n}^{(\mathbb{Z}_{q})}=H_{C}(n,\mathbb{Z}_{q},\boldsymbol{n/q})$ are given by $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})$ for all composition $\boldsymbol{r}=(r_{0},r_{1},\dots,r_{q-1})$.
Let $\boldsymbol{z}=(z_{i}:i\in\mathbb{Z}_{q})$ be a tuple of indeterminates, and let
$C(\boldsymbol{z})=(z_{j-i})_{i,j\in\mathbb{Z}_{q}}=\begin{pmatrix}z_{0}&z_{1}&\cdots&z_{q-2}&z_{q-1}\\ z_{q-1}&z_{0}&\cdots&z_{q-3}&z_{q-2}\\ \vdots&\vdots&&\vdots&\vdots\\ z_{2}&z_{3}&\cdots&z_{0}&z_{1}\\ z_{1}&z_{2}&\cdots&z_{q-1}&z_{0}\end{pmatrix}$
be the corresponding circulant matrix.
Observe that
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{r}}(\boldsymbol{n/q})=[\boldsymbol{z}^{\boldsymbol{r}}]\prod_{i\in\mathbb{Z}_{q}}\left(\sum_{j\in\mathbb{Z}_{q}}\zeta_{q}^{ij}z_{j}\right)^{n/q}=[\boldsymbol{z}^{\boldsymbol{r}}]\left(\prod_{i\in\mathbb{Z}_{q}}\sum_{j\in\mathbb{Z}_{q}}\zeta_{q}^{ij}z_{j}\right)^{n/q}=[\boldsymbol{z}^{\boldsymbol{r}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q},$
where the last equality follows from the well-known property of circulant matrices: the determinant equals the product of its eigenvalues $\sum_{j\in\mathbb{Z}_{q}}\zeta_{q}^{ij}z_{j}$ for $i=0,1,\dots,q-1$.
Therefore, using the reciprocal property, we obtain
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})=\frac{\binom{n}{\boldsymbol{n/q}}}{\binom{n}{\boldsymbol{r}}}\,[\boldsymbol{z}^{\boldsymbol{r}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}.$ (1)
Lemma 4.2.
For any composition $\boldsymbol{r}=(r_{0},\dots,r_{q-1})$, if $\sum_{i=0}^{q-1}ir_{i}\not\equiv 0\pmod{q}$, then $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})=0$.
Proof.
Let $e$ denote the all-one tuple in $\mathbb{Z}_{q}^{n}$, and let $a\in\mathbb{Z}_{q}^{n}$ be a tuple with $\operatorname{comp}(a)=\boldsymbol{r}$. Observe that
$e+S_{\boldsymbol{n/q}}=S_{\boldsymbol{n/q}}.$
Hence,
$\zeta_{q}^{a\cdot e}\cdot\sum_{x\in S_{\boldsymbol{n/q}}}\zeta_{q}^{a\cdot x}=\sum_{x\in S_{\boldsymbol{n/q}}}\zeta_{q}^{a\cdot(e+x)}=\sum_{x\in S_{\boldsymbol{n/q}}}\zeta_{q}^{a\cdot x}.$
If $a\cdot e=\sum_{i\in\mathbb{Z}_{q}}ir_{i}\not\equiv 0\pmod{q}$, then $\zeta_{q}^{a\cdot e}\neq 1$, which implies $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})=\sum_{x\in S_{\boldsymbol{n/q}}}\zeta_{q}^{a\cdot x}=0$. ∎
Lemma 4.3.
Let $\boldsymbol{r}=(r_{0},r_{1},\dots,r_{q-1})$ be a composition, and let $\boldsymbol{r}^{(1)}=(r_{1},r_{2},\dots,r_{q-1},r_{0})$ denote its left cyclic shift. Then
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r}^{(1)})=(-1)^{\frac{(q-1)n}{q}}K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r}).$
Proof.
Let $\boldsymbol{z}=(z_{i}:i\in\mathbb{Z}_{q})$ be a tuple of indeterminates, and let $\boldsymbol{z}^{(-1)}=(z_{q-1},z_{0},\dots,z_{q-2})$ denote its right cyclic shift. Then the circulant matrix $C(\boldsymbol{z})$ satisfies
$\det C(\boldsymbol{z}^{(-1)})=(-1)^{q-1}\det C(\boldsymbol{z}).$
Observe that $\boldsymbol{z}^{\boldsymbol{r}^{(1)}}=z_{0}^{r_{1}}z_{1}^{r_{2}}\cdots z_{q-2}^{r_{q-1}}z_{q-1}^{r_{0}}=z_{q-1}^{r_{0}}z_{0}^{r_{1}}z_{1}^{r_{2}}\cdots z_{q-2}^{r_{q-1}}=(\boldsymbol{z}^{(-1)})^{\boldsymbol{r}}$, we obtain
$\boldsymbol{z}^{\boldsymbol{r}^{(1)}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}=[(\boldsymbol{z}^{(-1)})^{\boldsymbol{r}}]\Big((-1)^{q-1}\det C(\boldsymbol{z}^{(-1)})\Big)^{n/q}=(-1)^{\frac{(q-1)n}{q}}[\boldsymbol{z}^{\boldsymbol{r}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}.$
Therefore,
$\begin{split}K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r}^{(1)})&=\frac{\binom{n}{\boldsymbol{n/q}}}{\binom{n}{\boldsymbol{r}^{(1)}}}\,[\boldsymbol{z}^{\boldsymbol{r}^{(1)}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}\\ &=\frac{\binom{n}{\boldsymbol{n/q}}}{\binom{n}{\boldsymbol{r}}}\,(-1)^{\frac{(q-1)n}{q}}[\boldsymbol{z}^{\boldsymbol{r}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}=(-1)^{\frac{(q-1)n}{q}}K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r}).\end{split}$
Lemma 4.4.
Let $q\geq 2$ be a positive integer. Let $n$ be divisible by $q$ such that $\frac{(q-1)n}{q}$ is an even integer. Then, there exists a constant $N(q)$ such that if $n\geq N(q)$, then the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$ is
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)=-\frac{\binom{n}{n/q,\ldots,n/q}}{n-1}.$
Proof.
Let $\boldsymbol{r}=(r_{0},r_{1},\ldots,r_{q-1})$ be a composition. By Lemma 4.3, since $\frac{(q-1)n}{q}$ is even, the value of generalized Krawtchouk polynomial $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}$ evaluated at $\boldsymbol{r}$ is invariant under cyclic shifts of $\boldsymbol{r}$. Therefore, without loss of generality, we may assume that $r_{0}\geq r_{i}$ for all $i=1,2,\ldots,q-1$, which in particular implies $r_{0}\geq n/q$.
Case 1. If $r_{0}=n$, then $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n,0,\dots,0)=\binom{n}{n/q,\dots,n/q}$ is the maximum eigenvalue.
Case 2. If $r_{0}=n-1$ and $r_{k}=1$ for some $k\neq 0$, then $\sum_{i\in\mathbb{Z}_{q}}ir_{i}=k\not\equiv 0\pmod{q}$, so by Lemma 4.2, $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})=0$.
Case 3. Suppose $r_{0}=n-2$. Then, by Lemma 4.2, the value $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})$ is nonzero only if there exists some $k\in\mathbb{Z}_{q}\setminus\{0\}$ such that either $k\neq-k$ and $r_{k}=r_{-k}=1$, or k=kk=-k and $r_{k}=2$. Therefore, in both subcases, we have
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})$ $=\frac{\binom{n}{\boldsymbol{n/q}}}{\binom{n}{\boldsymbol{r}}}[z_{0}^{n-2}z_{k}z_{-k}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}$
$=\frac{\binom{n}{\boldsymbol{n/q}}}{\binom{n}{\boldsymbol{r}}}\binom{n/q}{1}[z_{0}^{q-2}z_{k}z_{-k}]\det C(\boldsymbol{z})$
Since $\det C(\boldsymbol{z})$ is the determinant of the circulant matrix $C(\boldsymbol{z})$, by definition, we have
$\det C(\boldsymbol{z})=\sum_{\sigma\in\mathfrak{S}_{q}}\prod_{i=1}^{q}z_{\sigma(i)-i},$
where $\mathfrak{S}_{q}$ is the symmetric group of order $q$. The coefficient of $z_{0}^{q-2}$ in $\det C(\boldsymbol{z})$ only involve those $\sigma\in\mathfrak{S}_{q}$ that fix q2q-2 elements and transpose two remaining elements. We have
$[z_{0}^{q-2}]\det C(\boldsymbol{z})$ $=\sum_{\{k_{1},k_{2}\}\subseteq\mathbb{Z}_{q}}(-1)z_{k_{1}-k_{2}}z_{k_{2}-k_{1}}$
$=\begin{cases}-q(z_{1}z_{-1}+z_{2}z_{-2}+\cdots+z_{\frac{q-1}{2}}z_{-\frac{q-1}{2}}),&\text{ if $q$ is odd,}\\ -q(z_{1}z_{-1}+z_{2}z_{-2}+\cdots+z_{\frac{q}{2}-1}z_{-(\frac{q}{2}-1)})-\frac{q}{2}z_{\frac{q}{2}}^{2},&\text{ otherwise}.\end{cases}$
Therefore, in both subcases, we obtain $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})=-\frac{\binom{n}{n/q,\ldots,n/q}}{n-1}$.
Case 4. Suppose $n/q\leq r_{0}<n-\frac{q+3}{2}$. By Lemma 3.6, we have
$\left(\frac{K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})}{\binom{n}{n/q,\ldots,n/q}}\right)^{2}\leq\frac{q^{n}}{\binom{n}{n/q,\ldots,n/q}\binom{n}{r_{0},r_{1},\cdots,r_{q-1}}}$.
To show that $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)$ is the minimum eigenvalue, it suffices to show that
$\frac{q^{n}}{\binom{n}{n/q,\ldots,n/q}\binom{n}{r_{0},r_{1},\dots,r_{q-1}}}\leq\left(\frac{K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)}{\binom{n}{n/q,\ldots,n/q}}\right)^{2},$
since $K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)=-\frac{\binom{n}{n/q,\cdots,n/q}}{n-1}$, after ranging, it suffice to show that
$\binom{n}{r_{0},r_{1},\dots,r_{q-1}}\geq\frac{q^{n}(n-1)^{2}}{\binom{n}{n/q,\ldots,n/q}}$. (2)
Since $\frac{q+3}{2}<n/q\leq r_{0}<n-\frac{q+3}{2}$,
$\binom{n}{r_{0},r_{1},\ldots,r_{q-1}}\geq\binom{n}{r_{0},n-r_{0}}\geq\binom{n}{\lfloor\frac{q+3}{2}\rfloor+1}\gtrsim_{q}n^{\lfloor\frac{q+3}{2}\rfloor+1}.$
On the other hand, applying Stirling’s approximation, we have
$\frac{q^{n}(n-1)^{2}}{\binom{n}{n/q,\ldots,n/q}}\leq\frac{(n-1)^{2}\left(\sqrt{2\pi n/q}\right)^{q}e^{\frac{q^{2}}{12n}}}{\sqrt{2\pi n}}\lesssim_{q}n^{\frac{q+3}{2}}.$
Therefore, (2) is satisfied when $n$ is sufficiently large.
Case 5. Suppose $n-\frac{q+3}{2}\leq r_{0}\leq n-3$. Let $3\leq s\leq\lfloor\frac{q+3}{2}\rfloor$ and $r_{0}=n-s$. Let $\operatorname{per}C(\boldsymbol{z})$ denote the permanent of $C(\boldsymbol{z})$. Then, it is clear that
$\Big|[z_{0}^{r_{0}}z_{1}^{r_{1}}\cdots z_{q-1}^{r_{q-1}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}\Big|\leq[z_{0}^{r_{0}}z_{1}^{r_{1}}\cdots z_{q-1}^{r_{q-1}}]\Big(\operatorname{per}C(\boldsymbol{z})\Big)^{n/q}\leq[z_{0}^{r_{0}}]\Big(\operatorname{per}C(z_{0},1,\dots,1)\Big)^{n/q}.$ (3)
Since the coefficient of $z_{0}^{q-1}$ in the expansion of
[…]
must be zero. Therefore, we can write
$\operatorname{per}C(z_{0},1,\ldots,1)=a_{q}z_{0}^{q}+a_{q-1}z_{0}^{q-1}+a_{q-2}z_{0}^{q-2}+\cdots+a_{0},$
for some nonnegative integers $a_{q}=1,a_{q-1}=0$, and $a_{i}\leq q!$, $i=0,1,\ldots,q-2$. Then
$[z_{0}^{r_{0}}]\Big(\operatorname{per}C(z_{0},1,\ldots,1)\Big)^{n/q}=[z_{0}^{n-s}]\left(\sum_{i=0}^{q}a_{i}z_{0}^{i}\right)^{n/q}=\sum_{\begin{subarray}{c}k_{1}+k_{2}+\cdots+k_{n/q}=n-s\\ \text{each }k_{j}\neq q-1\end{subarray}}a_{k_{1}}a_{k_{2}}\cdots a_{k_{n/q}}.$ (4)
Let $\boldsymbol{k}=(k_{1},k_{2},\ldots,k_{n/q})$ denote a feasible solution to $\sum_{j=1}^{n/q}k_{j}=n-s$ with $k_{j}\neq q-1$ for all $j\in[n/q]$, and let $N(\boldsymbol{k})$ be the number of $j\in[n/q]$ such that $k_{j}\leq q-2$. Since $k_{j}\neq q-1$ for all $j\in[n/q]$, we have $n-s=\sum_{j=1}^{n/q}k_{j}\leq q-2N(\boldsymbol{k})$, implying that $N(\boldsymbol{k})\leq\frac{s}{2}$. Therefore,
$\sum_{\begin{subarray}{c}k_{1}+k_{2}+\cdots+k_{n/q}=n-i\\ \text{each }k_{j}\neq q-1\end{subarray}}a_{k_{1}}a_{k_{2}}\cdots a_{k_{n/q}}$ $=\sum_{\ell=0}^{\lfloor\frac{s}{2}\rfloor}\sum_{N(\boldsymbol{k})=\ell}a_{k_{1}}a_{k_{2}}\cdots a_{k_{n/q}}$ (5)
$\leq\sum_{\ell=0}^{\lfloor\frac{s}{2}\rfloor}\sum_{N(\boldsymbol{k})=\ell}(q!)^{\ell}\leq\sum_{\ell=0}^{\lfloor\frac{s}{2}\rfloor}\binom{n/q}{\ell}(q-1)^{\ell}(q!)^{\ell}\lesssim_{q}n^{\frac{s}{2}}.$
Combining (3),(4) and (5), we obtain
$\Big|[z_{0}^{r_{0}}z_{1}^{r_{1}}\cdots z_{q-1}^{r_{q-1}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}\Big|\lesssim_{q}n^{\frac{s}{2}}.$
Therefore,
$\left|\frac{K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})}{\binom{n}{n/q,\dots,n/q}}\right|=\frac{\Big|[z_{0}^{r_{0}}z_{1}^{r_{1}}\cdots z_{q-1}^{r_{q-1}}]\Big(\det C(\boldsymbol{z})\Big)^{n/q}\Big|}{\binom{n}{n-s,r_{1},\dots,r_{q-1}}}\lesssim_{q}\frac{n^{\frac{s}{2}}}{\binom{n}{s}}\lesssim_{q}n^{-\frac{s}{2}},$
which decreases faster than
$\left|\frac{K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)}{\binom{n}{n/q,\dots,n/q}}\right|=\frac{1}{n-1}.$
Hence $\left|K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(\boldsymbol{r})\right|\leq\left|K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)\right|$ for sufficiently large $n$. ∎
4.2 Minimum eigenvalue of $\varOmega_{n}^{(\mathbb{F}_{q})}$
Here, we present an argument of algebraic flavor to determine the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{F}_{q})}$. We first consider the case $n=q$, and then lift it to the prime power case where $n$ is divisible by $q$.
Lemma 4.5.
For any prime power $q$, the minimum eigenvalue of $\varOmega_{q}^{(\mathbb{F}_{q})}$ is $\frac{q!}{q-1}$.
Proof.
Recall that $\varOmega_{q}^{(\mathbb{F}_{q})}=\mathrm{Cay}(\mathbb{F}_{q}^{q},S_{\boldsymbol{1}})$, where $S_{\boldsymbol{1}}$ is the set of vectors in $\mathbb{F}_{q}^{q}$ having composition $\boldsymbol{1}=(1,1,\ldots,1)$.
Suppose that $q$ is a power of a prime $p$. By Theorem 2.3, the eigenvalues of $\varOmega_{q}^{(\mathbb{F}_{q})}$ are
$\mu_{a}=\sum_{s\in S_{\boldsymbol{1}}}\zeta_{p}^{\operatorname{Tr}(s\cdot a)},\quad a\in\mathbb{F}_{q}^{q}.$
Observe that for each $x\in\mathbb{F}_{q}^{\times}$, we have $xS_{\boldsymbol{1}}=\{xs:s\in S_{\boldsymbol{1}}\}=S_{\boldsymbol{1}}$. Therefore,
$\begin{split}\mu_{a}=\sum_{s\in S_{\boldsymbol{1}}}\zeta_{p}^{\operatorname{Tr}(a\cdot s)}&=\frac{1}{q-1}\sum_{x\in\mathbb{F}_{q}^{\times}}\sum_{s\in S_{\boldsymbol{1}}}\zeta_{p}^{\operatorname{Tr}(a\cdot(xs))}\\ &=\frac{1}{q-1}\sum_{s\in S_{\boldsymbol{1}}}\sum_{x\in\mathbb{F}_{q}^{\times}}\zeta_{p}^{\operatorname{Tr}(x(a\cdot s))}\\ &=\frac{1}{q-1}\left(\sum_{s\in S_{\boldsymbol{1}}:a\cdot s=0}(q-1)+\sum_{s\in S_{\boldsymbol{1}}:a\cdot s\neq 0}(-1)\right)\\ &=\frac{1}{q-1}\left(\sum_{s\in S_{\boldsymbol{1}}:a\cdot s=0}q-\sum_{s\in S_{\boldsymbol{1}}}1\right)\geq-\frac{|S_{\boldsymbol{1}}|}{q-1},\end{split}$
where the last inequality becomes equality if $a=(1,-1,0,\ldots,0)$, since there is no $s\in S$ that satisfies $a\cdot s=0$. Thus, the minimum eigenvalue of $\varOmega_{q}^{(\mathbb{F}_{q})}$ is $\frac{q!}{q-1}$. ∎
Lemma 4.6.
Let $q$ and $n$ be prime powers with $q\mid n$. Then, the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{F}_{q})}$ is $\frac{\binom{n}{n/q,\dots,n/q}}{n-1}$.
Proof.
Recall that $\varOmega_{n}^{(\mathbb{F}_{n})}=\mathrm{Cay}(\mathbb{F}_{n}^{n},S_{\boldsymbol{1}})$ and $\varOmega_{n}^{(\mathbb{F}_{q})}=\mathrm{Cay}(\mathbb{F}_{q}^{n},S_{\boldsymbol{n/q}})$.
Let $\theta_{0}:\mathbb{F}_{n}\rightarrow\mathbb{F}_{q}$ be any surjective linear map over the prime field, and let $\theta:\mathbb{F}_{n}^{n}\rightarrow\mathbb{F}_{q}^{n}$ be the map induced by $\theta_{0}$, defined componentwise by
$\theta(a_{1},\dots,a_{n})=(\theta_{0}(a_{1}),\dots,\theta_{0}(a_{n})).$
For a character $\psi$ of $\mathbb{F}_{q}^{n}$, we write $\psi(S):=\sum_{s\in S}\psi(s)$ for brevity. Since every element of $S_{\boldsymbol{n/q}}$ has the same number of preimages in $S_{\boldsymbol{1}}$, it follows that
$(\psi\circ\theta)(S_{\boldsymbol{1}})=\frac{|S_{\boldsymbol{1}}|}{|S_{\boldsymbol{n/q}}|}\,\psi(S_{\boldsymbol{n/q}}).$
Since $\psi\circ\theta$ is a character of $\mathbb{F}_{n}^{n}$, $(\psi\circ\theta)(S_{\boldsymbol{1}})$ is an eigenvalue of $\varOmega_{n}^{(\mathbb{F}_{n})}$. By Lemma 4.5, we have
$(\psi\circ\theta)(S_{\boldsymbol{1}})\geq-\frac{|S_{\boldsymbol{1}}|}{n-1},$
which immediately implies
$\psi(S_{\boldsymbol{n/q}})\geq-\frac{|S_{\boldsymbol{n/q}}|}{n-1}.$
Thus, every eigenvalue of $\varOmega_{n}^{(\mathbb{F}_{q})}$ is lower bounded by $\frac{|S_{\boldsymbol{n/q}}|}{n-1}$.
To see that this bound is attained, consider $a=(1,-1,0,\dots,0)\in\mathbb{F}_{q}^{n}$ and the character $\psi_{a}$ defined by $\psi_{a}(x)=\zeta_{p}^{\mathrm{Tr}(a\cdot x)}$.
Observe that for each $x\in\mathbb{F}_{q}^{\times}$, $xS_{\boldsymbol{n/q}}=S_{\boldsymbol{n/q}}$. In other words, there is a natural action of the multiplicative group $\mathbb{F}_{q}^{\times}$ on $S_{\boldsymbol{n/q}}$ by scalar multiplication. Denote the set of orbits by $S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}$; each orbit contains exactly q1q-1 elements, and these orbits partition $S_{\boldsymbol{n/q}}$.
Then we can write
$\psi_{a}(S_{\boldsymbol{n/q}})$ $=\sum_{x\in S_{\boldsymbol{n/q}}}\zeta_{p}^{\mathrm{Tr}(x_{1}-x_{2})}$
$=\sum_{[x]\in S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}}\sum_{\alpha\in\mathbb{F}_{q}^{\times}}\zeta_{p}^{\mathrm{Tr}(\alpha(x_{1}-x_{2}))}$
$=\sum_{[x]\in S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}:x_{1}=x_{2}}(q-1)+\sum_{[x]\in S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}:x_{1}\neq x_{2}}(-1)$
$=\sum_{[x]\in S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}:x_{1}=x_{2}}q-\sum_{[x]\in S_{\boldsymbol{n/q}}/\mathbb{F}_{q}^{\times}}1$
$=\frac{q-1}{q}\bigl|\{x\in S_{\boldsymbol{n/q}}:x_{1}=x_{2}\}\bigr|-\frac{|S_{\boldsymbol{n/q}}|}{q-1}$
$=\frac{q-1}{q}\binom{q}{1}\binom{n-2}{n/q-2,n/q,\dots,n/q}-\frac{|S_{\boldsymbol{n/q}}|}{q-1}$
$=-\frac{|S_{\boldsymbol{n/q}}|}{n-1}.$
Hence, the lower bound is attained, completing the proof. ∎
Proof of Theorem 1.5.
Theorem 1.5 follows from Lemmas 2.1, 4.1, 2.2, 4.4 and 4.6. ∎
5 Concluding remarks
We investigate the quantum chromatic number of Hamming and generalized Hadamard graphs. Several interesting open questions arise from our work.
Orthogonal representations for $H(n,q,d)$ with $d<\frac{(q-1)n}{q}$.
We develop a linear programming approach to construct modules-one orthogonal representations for Hamming graphs and establish upper bounds on $\xi^{\prime}(H(n,q,d))$ for $d\leq\frac{(q-1)n}{q}$. However, the bound in Corollary 3.4 remains exponentially large.
Question 5.1.
For $d=\delta n$ with $0<\delta<\frac{q-1}{q}$, and with $d$ even when $q=2$, is it always true that
$\xi^{\prime}(H(n,q,d))\leq\operatorname{poly}(n),$
or does there exist some $d$ such that
$\xi^{\prime}(H(n,q,d))\geq\operatorname{exp}(n)?$
Exact value of $\chi_{Q}(H(n,q,\frac{(q-1)n}{q}))$ for $q\geq 3$.
We show that
$(q-1)(n-1)+1\leq\chi_{Q}\big(H(n,q,\tfrac{(q-1)n}{q})\big)\leq(q-1)n,$
but a gap of (q2)(q-2) remains between the upper and lower bounds.
Question 5.2.
How can this gap be closed?
Determining the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$ for all $n$.
We determine the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$ for sufficiently large $n$, and we conjecture that it remains the same for all feasible $n$. Formally, we state the conjecture as follows:
Conjecture 5.3.
Let $q\geq 2$ be a positive integer, and let $n$ be divisible by $q$ such that $\frac{(q-1)n}{q}$ is an even integer. Then the minimum eigenvalue of $\varOmega_{n}^{(\mathbb{Z}_{q})}$ is
$K^{(\mathbb{Z}_{q})}_{\boldsymbol{n/q}}(n-2,1,0,\dots,0,1)=-\frac{\binom{n}{n/q,\ldots,n/q}}{n-1}.$